# Meta-Reasoning Improves Tool Use in Large Language Models

### v1

@@ Abstract
External tools help large language models (LLMs) succeed at tasks where they would otherwise typically fail. In existing frameworks, LLMs learn tool use either by in-context demonstrations or via full model fine-tuning on annotated data. As these approaches do not easily scale, a recent trend is to abandon them in favor of lightweight, parameter-efficient tuning paradigms. These methods allow quickly alternating between the frozen LLM and its specialised fine-tuned version, by switching on or off a handful of additional custom parameters. Hence, we postulate that the generalization ability of the frozen model can be leveraged to improve tool selection. We present Tool selection via meta-reasoning (Tecton), a two-phase system that first reasons over a task using a custom fine-tuned LM head and outputs candidate tools. Then, with the custom head disabled, it meta-reasons (i.e., it reasons over the previous reasoning process) to make a final choice. We show that Tecton results in substantial gains – both in-distribution and out-of-distribution – on a range of math reasoning datasets.
@@ 1 Introduction
Augmentation with external tools has proven effective at boosting the performance of large language models in knowledge-intensive tasks such as QA and math problem-solving Hao et al. (2023); Paranjape et al. (2023); Parisi et al. (2022); Schick et al. (2023). Tools are self-contained programs or APIs whose execution timing and arguments are determined by the language model. To teach a model how to use tools, previous works adopt one of two main strategies: (1) tool demonstrations via in-context learning Gao et al. (2023); Gupta and Kembhavi (2023); Hsieh et al. (2023); Surís et al. (2023), or (2) fine-tuning on a dataset where text samples are interleaved with tool annotations Parisi et al. (2022); Schick et al. (2023); Tang et al. (2023). Hao et al. (2023) note that a major downside of in-context demonstrations is the finite size of the context itself, preventing the system from being able to scale up to a very large number of tools. The issue of scalability is not just theoretical: according to Patil et al. (2023), a general-purpose LLM may need access to thousands of tools to be able to carry out complex and diverse tasks (e.g., booking holidays or organizing conferences) from natural language instructions. In contrast, fine-tuned models are able to learn thousands of tools from annotated data Patil et al. (2023). However, they are bound to the set of tools learned during training: adding further tools requires a new round of fine-tuning at significant computational cost Hao et al. (2023).
Recent works adopt parameter-efficient learning to reduce the cost of fine-tuning, thus facilitating potential future extensions of the tool set. Hao et al. (2023) develop ToolkenGPT by augmenting the output matrix of an LLM with additional special tokens, each representing a tool. Only the additional tokens are tuned while the rest of the weights remain frozen. Similarly, Wang et al. (2024) train tokens representing math operations to aid each reasoning step, using LoRA Hu et al. (2022) to minimize the number of additional parameters. Qiao et al. (2024) implement a two-stage system called Trice, where instruction tuning is followed by reinforcement learning from environmental feedback (RLEF), leveraging LoRA at both stages.
@@
An advantage of parameter-efficient paradigms is that they preserve most of the LLM’s core capabilities acquired at pre-training Ding et al. (2023); Han et al. (2024). These methods only tune a handful of additional parameters for a specific task, which can be selectively disabled to reinstate the original model. We argue that previous works exploring these techniques for tool generation do not fully leverage the generalization power of the underlying LLM. To address this issue, we propose a two-phase framework for Tool selection via meta-reasoning (Tecton). In the reasoning phase, Tecton investigates a task and outputs candidate tools with the aid of a custom-tuned language modeling head. Then, in the meta-reasoning phase, it uses the frozen LLM to re-examine these candidates and make a final decision. This differs from previous works where the choice of tool is solely determined by its probability at decoding time Hao et al. (2023); Wang et al. (2024); Schick et al. (2023). Fig. 1 illustrates the system. We train and evaluate Tecton on math reasoning datasets, following established works on LLM tool calling Chen et al. (2024); Das et al. (2024); Gou et al. (2024). It is worth noting that math reasoning problems are highly challenging, since they require chains of multiple tool calls with errors that compound. This is evidenced by existing tool-augmented LLMs, which achieve the lowest performance on math reasoning when evaluated on multiple tasks Hao et al. (2023); Schick et al. (2023).
In summary, our main contributions are:
@@
  • We introduce Tecton, a novel two-phase framework that combines a custom fine-tuned head with a frozen LLM to improve tool use in math reasoning tasks (Section 2).
@@
  • We show that Tecton outperforms strong baselines both on in-distribution data and on unseen benchmarks (Section 3).
@@ 2 Method
We augment the language modeling head of a base model with additional token embeddings $T$ to represent math operations, and train them via a standard language modeling objective. Existing works have shown the effectiveness of tuning additional tokens for math tasks Hao et al. (2023); Wang et al. (2024) and general reasoning Goyal et al. (2024); Herel and Mikolov (2024). To take full advantage of both the specialized tokens and the generalist frozen LLM, we design a two-phase framework. The rest of the section describes this in detail.
@@ 2.1 Reasoning Phase
In the reasoning phase, we ask the LLM to solve a math problem step by step. Rather than greedily generate the most likely tool as in previous works Hao et al. (2023); Wang et al. (2024), we leverage the augmented LM head to collect a set of candidate tools. To this end, we experiment with both temperature sampling and greedy decoding of multiple tokens (see Appendix B). We find that looking at the top $k$ most likely tokens at every position in the sequence is the most promising approach to gather a diverse yet relevant pool of candidates. Hence, we greedily generate text to solve each math problem line by line, gathering tools from the top $k$ tokens at each decoding step. For each line of text, the multiset $C$ of candidate tools is thus given by
@@
where $T$ is the set of available tool tokens, and $W$ is the matrix resulting from decoding $k$ top-probability tokens at each of the $l$ token positions in the line of text. We set $k=5$ as a tradeoff between search space size and computation cost. We then prompt the LLM to produce arguments for each candidate tool given its previous context. Identical tools with the same arguments are dropped from the pool. Finally, we pass each candidate tool and arguments into the Python interpreter, and keep those that are successfully evaluated. Once a line of text toward the solution has been processed in this way, we move onto the meta-reasoning phase.
@@ 2.2 Meta-Reasoning Phase
In this phase, we let the LLM analyze its previous reasoning process and choose among the candidate tools. To this end, we use the frozen model without custom tokens. Pre-trained language models have been used for self-evaluation in previous literature Alazraki et al. (2023); Shinn et al. (2023); Yao et al. (2023a). We experiment with two ways of eliciting meta-reasoning:
  1. 1.
    We join each candidate tool with the previous text context, and present these as options for the LLM to score. We prefix each option with an uppercase letter label and select as the answer continuation the option whose label is assigned highest probability by the model, i.e., $\text{arg max}_{t_{i}\in\mathcal{V}_{\text{sub}}}\,p(t_{i}\mid t_{{}<i}\text{ {with} }t_{{}<i}\in\mathcal{V})$, where $\mathcal{V}_{\text{sub}}$ denotes a subset of the vocabulary containing only the uppercase letter tokens that are in the label set. In this setup, which we call Tecton-score, we limit the number of candidates to a maximum of four.
  2. 2.
    We pass the candidate tools as hints and ask the model to generate an appropriate continuation of the answer. Here, the hints serve as mere guidance for the LLM (i.e., the model could choose to ignore all candidates and generate something different). In this version of the system, which we call Tecton-generate, the model benefits from dynamically retrieved few-shot exemplars demonstrating the tools in the candidate set.
In both setups, the model additionally observes the original problem description to aid its decision.
@@ 3.1 Experimental Setup
Our system is model-agnostic and can be applied to any open-weights LLM. Here, we use Llama 3 8B Instruct Dubey et al. (2024) (from here on referred to as Llama 3) as the base model in all experiments. Implementation details and hyperparameters are given in Appendix A.
@@ 3.2 Datasets
We train and evaluate Tecton on GSM8K-XL Cobbe et al. (2021); Hao et al. (2023) and FuncQA Hao et al. (2023). The test set of the latter is comprised of two distinct subsets: a ‘one-hop’ corpus containing problems solvable with one single operation (FuncQA-OH), and a ‘multi-hop’ one requiring multiple operations (FuncQA-MH). For GSM8K-XL we tune four additional tokens corresponding to the four basic operations, and extend these to 13 in the case of FuncQA. The complete set of tools for each dataset is shown in Appendix A. Additionally, we evaluate on out-of-distribution datasets that were not observed during fine-tuning. For this purpose we choose a range of math reasoning datasets: ASDiv Miao et al. (2020), MAWPS Koncel-Kedziorski et al. (2016) and SVAMP Patel et al. (2021). These were found by Ott et al. (2023) to be OOD with respect to GSM8K: not only do they display minimal n-gram overlap, but they also require reasoning chains of different average lengths from GSM8K to be solved. These datasets have been used to evaluate LLMs in previous works, both with and without the aid of tools Kojima et al. (2022); Schick et al. (2023). Following a strategy similar to the one used by Hao et al. (2023) for constructing GSM8K-XL, we magnify the numbers in these datasets to make them challenging for current LLMs (the enhancement process is described in Appendix D). We thus obtain ASDiv-XL, MAWPS-XL, and SVAMP-XL. We use these datasets to test models tuned on GSM8K-XL.
@@ 3.3 Baselines
We implement recent tool-augmented models as baselines: Trice Qiao et al. (2024) and ToolkenGPT Hao et al. (2023). These share a parameter-efficient approach with Tecton. Despite their relatively low computational cost, they have been shown to outperform strong systems: Trice paired with Alpaca Taori et al. (2023), ChatGLM Zeng et al. (2024) and Vicuna Zheng et al. (2023) surpasses the much larger GPT-3.5 as well as tool learning via supervised fine-tuning. In addition to outperforming GPT-3.5, ToolkenGPT paired with LLaMA Touvron et al. (2023) is more accurate than ReAct Yao et al. (2023b). It should also be noted that LLMs are increasingly able to solve math problems and perform difficult arithmetic without tools, as shown by the relatively high accuracy ($79.6\%$) obtained by Llama 3 on the non-enhanced version of GSM8K Dubey et al. (2024). Therefore, we additionally compare against a vanilla version of Llama 3 as well as Chain-of-Thought (CoT) prompting Wei et al. (2022) with exemplars extracted from the training set.
@@ 3.4 Results
In-distribution performance  On GSM8K-XL, our best implementation scores $7.2$ percentage points above Trice and $2.3$ above ToolkenGPT. When evaluated on in-distribution data, Tecton’s most significant gains are on FuncQA: the generate setting doubles the performance of ToolkenGPT in the challenging multi-hop task, reaching $20.6\%$ accuracy. For context, Llama 3 can only solve $2.9\%$ of the dataset and CoT only raises performance to $5.9\%$.
Out-of-distribution performance  On OOD datasets, Tecton’s advantage is substantial: on average, Tecton-generate gains $8.5$ percentage points over ToolkenGPT and $8.8$ over Trice. Tecton-score achieves $6.2$ and $6.6$ above the two baselines, respectively. This highlights the adaptability of the method to unseen data.
@@ Acknowledgments
Appendix A Model Implementation and Training
@@
A.1 Tools
GSM8K-XL and FuncQA Hao et al. (2023) are annotated with four and 13 tools respectively, each representing a math operation. We illustrate these in Table 3. Tools are trained as additional tokens added to the standard language modeling head of Llama 3, which comprises 128,256 token representations. Therefore, we extend these representations to 128,269 in the case of FuncQA and 128,260 for GSM8K-XL.
A.2 Details of the Training Process
The LM head of Tecton consists of the standard head of Llama 3 8B Instruct Dubey et al. (2024) concatenated with an additional linear layer of size $embedding\_dimension\times number\_of\_tools$. The tool token embeddings are randomly initialized and trained on math reasoning QA pairs from GSM8K-XL and FuncQA, following a standard language modeling objective. Both datasets are annotated with the token positions at which each tool should be generated. Note that the original datasets made available by Hao et al. (2023) are annotated according to SentencePiece tokenization Kudo and Richardson (2018). We edit the annotations to be compatible with Llama 3’s Byte-Pair Encoding tokenizer Sennrich et al. (2016). We tune two distinct sets of special tokens, one on each dataset, as each requires a different set of tools as shown in Table 3. We train at half precision (FP16) for ten epochs, using learning rates {$1$e44-4- 4, $1$e33-3- 3}, batch size 1, and saving checkpoints at each epoch. We select the best checkpoint by measuring performance on a validation set.
@@ A.2 Details of the Training Process
Table 4: Sizes of our training and testing datasets. Note that we use ASDiv-XL, MAWPS-XL and SVAMP-XL only for testing. For FuncQA the reported test data size includes both splits – FuncQA-OH and FuncQA-MH.
@@
A.3 Inference-time Hyperparameters

### v2

@@ Abstract
External tools help large language models succeed at tasks where they would otherwise typically fail. In existing frameworks, choosing tools at test time relies on naive greedy decoding, regardless of whether the model has been fine-tuned on tool-annotated data or prompted with in-context examples. In contrast, we find that gathering and choosing among a suitable set of candidate tools has greater potential to lead to an optimal selection. We present Tool selection via meta-reasoning (Tecton), a two-phase system that first reasons over a task and outputs candidate tools using a custom fine-tuned language modelling head. Then, with the custom head disabled, it meta-reasons (i.e., it reasons over the previous reasoning process) to make a final choice. We show that Tecton results in substantial gains—both in-distribution and out-of-distribution—on a range of math reasoning datasets.
@@ 1 Introduction
Augmentation with external tools has proven effective at boosting the performance of large language models (LLMs) in knowledge-intensive tasks such as QA and math problem-solving Hao et al. (2023); Paranjape et al. (2023); Parisi et al. (2022); Schick et al. (2023). Tools are self-contained programs or APIs which the model can execute with chosen arguments as needed. To teach an LLM how to use tools, previous work adopts one of three main strategies: (1) tool demonstrations via in-context learning (ICL) Gao et al. (2023); Gupta and Kembhavi (2023); Hsieh et al. (2023); Surís et al. (2023), (2) full model fine-tuning on a dataset where text samples are interleaved with tool annotations Parisi et al. (2022); Schick et al. (2023); Tang et al. (2023); Patil et al. (2023), or (3) parameter-efficient fine-tuning (PEFT) on tool annotated data Hao et al. (2023); Qiao et al. (2024); Wang et al. (2024). Similarly to full fine-tuning, PEFT methods can teach LLMs a very large number of tools, while ICL is limited by the fixed size of the context window Hao et al. (2023); Patil et al. (2023). Although the fine-tuning paradigm in general binds the model to the set of tools learned during training, parameter-efficient learning reduces the cost of tuning, thus facilitating potential future extensions of the tool set. In contrast, adding further tools via a new round of full-model tuning incurs a significant computational cost Hao et al. (2023). We note that a further advantage of PEFT is that it tunes a handful of additional task-specific parameters, which can be selectively disabled to reinstate the frozen model and its original capabilities Ding et al. (2023); Han et al. (2024).
@@
In existing literature, inference-time tool selection is made by greedily decoding the most likely tool Gao et al. (2023); Hao et al. (2023); Schick et al. (2023); Wang et al. (2024), regardless of whether the model has been fully fine-tuned, PEFT-tuned, or prompted in-context. In this work, we revisit the paradigm that selects tools solely based on their probability at decoding time. We propose an alternative, novel framework for Tool selection via meta-reasoning (Tecton) that gathers and chooses among a suitable set of candidate tools. We train a parameter-efficient language modelling head on tool-annotated data, similar to Hao et al. (2023), which can be switched on or off as needed. Tecton thus comprises two distinct phases: in the reasoning phase, it investigates a task and outputs candidate tools with the aid of the tuned LM head. Then, in the meta-reasoning phase, it uses the frozen LLM to re-examine the candidates and make a final decision. Fig. 1 illustrates the framework. We train and evaluate Tecton on math reasoning datasets, following established work on LLM tool calling Chen et al. (2024); Das et al. (2024); Gou et al. (2024). Among tasks that benefit from tools, math reasoning is particularly challenging, since it requires chains of multiple tool calls with errors that compound. This is evidenced by existing tool-augmented LLMs, which achieve the lowest performance on math reasoning when evaluated on multiple tasks Hao et al. (2023); Schick et al. (2023). In summary, our main contributions are:
@@
  • We introduce Tecton, a novel two-phase framework that combines a custom fine-tuned head with a frozen LLM to improve tool use (Section 2).
@@
  • We show that Tecton outperforms strong baselines in math reasoning tasks, both on in-distribution data and on unseen benchmarks (Section 3).
@@ 2 Method
2.1 Preliminaries
We augment the language modelling head of a base model with additional token embeddings $T$ to represent math operations, and train them via a standard language modelling objective. Once trained, $T$ comprises the tools available to the LLM for solving math problems. Prior work has shown the effectiveness of tuning additional tokens for both math tasks Hao et al. (2023); Wang et al. (2024) and general reasoning Goyal et al. (2024); Herel and Mikolov (2024).
Our preliminary experiments show that in cases where the LLM has failed to generate the correct tool by greedy sampling, this can usually be found among tokens that have only slightly lower probabilities (Figure 3). Over-sampling an appropriate set of candidate tools may thus be a better strategy than greedily decoding the most likely tool, provided this is combined with a reliable method for choosing among the candidates. To this end, we design a two-phase framework that leverages both the specialised, augmented LM head (reasoning phase) and the underlying generalist LLM (meta-reasoning phase). The rest of the section describes this in detail.
@@ 2.3 Meta-Reasoning Phase
In this phase, we disable the custom-tuned head and let the underlying LLM analyse its previous reasoning process to choose among the candidate tools. Frozen LLMs have been used to self-evaluate and meta-reason over previous answers in existing literature Alazraki et al. (2023); Shinn et al. (2023); Yao et al. (2023a); Zeng et al. (2024b). We experiment with two ways of eliciting meta-reasoning: Tecton-score and Tecton-generate.
Tecton-score.
We join each candidate tool with the previous context, and present these as options for the LLM to score. We prefix each option with an uppercase letter label and select as the answer continuation the option whose label is assigned highest probability by the model, i.e., $\text{arg max}_{t_{i}\in\mathcal{V}_{\text{sub}}}\,p(t_{i}\mid t_{{}<i}\text{ {with} }t_{{}<i}\in\mathcal{V})$, where $\mathcal{V}_{\text{sub}}$ denotes a subset of the vocabulary containing only the uppercase letter tokens that are in the label set. In this setup, we limit the number of candidates to a maximum of four.
Tecton-generate.
We pass the candidate tools as hints and ask the model to generate an appropriate continuation of the answer. Here, the hints serve as mere guidance for the LLM (i.e., the model could choose to ignore all candidates and generate something different). In this version of the system, the model benefits from dynamically retrieved few-shot exemplars demonstrating the tools in the candidate set.
@@ 3.1 Experimental Setup
Our system is model-agnostic and can be applied to any open-weights LLM. Here, we use Llama 3 8B Instruct Dubey et al. (2024) (henceforth referred to as Llama 3) as the base model in all experiments. Implementation details and hyperparameters are given in Appendix B.
@@ 3.2 Datasets
We train and evaluate Tecton on GSM8K-XL Cobbe et al. (2021); Hao et al. (2023) and FuncQA Hao et al. (2023). The test set of the latter is comprised of two distinct subsets: a ‘one-hop’ corpus containing problems solvable with one single operation (FuncQA-OH), and a ‘multi-hop’ one requiring multiple operations (FuncQA-MH). For GSM8K-XL we tune four additional tokens corresponding to the four basic operations, and extend these to 13 in the case of FuncQA. The complete set of tools for each dataset is shown in Appendix B. Additionally, we evaluate on out-of-distribution datasets that were not observed during fine-tuning. For this purpose we choose a range of math reasoning datasets: ASDiv Miao et al. (2020), MAWPS Koncel-Kedziorski et al. (2016) and SVAMP Patel et al. (2021). These were found by Ott et al. (2023) to be OOD with respect to GSM8K: not only do they display minimal n-gram overlap, but they also require reasoning chains of different average lengths from GSM8K to be solved. These datasets have been used to evaluate LLMs in previous works, both with and without the aid of tools Kojima et al. (2022); Schick et al. (2023). Following a strategy similar to the one used by Hao et al. (2023) for constructing GSM8K-XL, we magnify the numbers in these datasets to make them challenging for current LLMs (the enhancement process is described in Appendix D). We thus obtain ASDiv-XL, MAWPS-XL, and SVAMP-XL. We use these datasets to test models tuned on GSM8K-XL.
@@ 3.3 Baselines
We implement recent tool-augmented models as baselines: Trice Qiao et al. (2024) and ToolkenGPT Hao et al. (2023). These share a parameter-efficient approach with Tecton (see Appendix C for details). Despite their relatively low computational cost, they have been shown to outperform strong systems: Trice paired with Alpaca Taori et al. (2023), ChatGLM Zeng et al. (2024a) and Vicuna Zheng et al. (2023) surpasses the much larger GPT-3.5 as well as tool learning via supervised fine-tuning. In addition to outperforming GPT-3.5, ToolkenGPT paired with LLaMA Touvron et al. (2023) is more accurate than ReAct Yao et al. (2023b). It should also be noted that LLMs are increasingly able to solve math problems and perform difficult arithmetic without tools, as shown by the high accuracy ($79.6\%$) achieved by Llama 3 on the non-enhanced version of GSM8K Dubey et al. (2024). Hence we additionally compare against a vanilla version of Llama 3 as well as Chain-of-Thought (CoT) prompting Wei et al. (2022) with exemplars extracted from the train set.
@@ 3.4 Results
In-distribution performance.
On GSM8K-XL, our best implementation scores $7.2$ percentage points above Trice and $2.3$ above ToolkenGPT. When evaluated on in-distribution data, Tecton’s most significant gains are on FuncQA: the generate setting doubles the performance of ToolkenGPT in the challenging multi-hop task, reaching $20.6\%$ accuracy. For context, Llama 3 can only solve $2.9\%$ of the dataset and CoT only raises performance to $5.9\%$.
Out-of-distribution performance.
On OOD datasets, Tecton’s advantage is substantial: on average, Tecton-generate gains $8.5$ percentage points over ToolkenGPT and $8.8$ over Trice. Tecton-score achieves $6.2$ and $6.6$ above the two baselines, respectively. This highlights the adaptability of the method to unseen data.
@@ Acknowledgments
Appendix A Preliminary Experiments
A.1 Top-k Decoding
We run ToolkenGPT (in its original implementation based on the first version of LLaMA) on the validation set of FuncQA. We find that in samples where the system has generated an incorrect tool by greedy sampling, the correct one—complete with correct arguments—is among the top five most likely tokens in over $60\%$ of cases. These include samples where the correct tool can be found by searching the top tokens at a different position from the one that has produced the incorrect tool. Overall, the system decodes the correct tool (regardless of the arguments it generates for it) in 76.9% of samples; this rises by over $10\%$ to $87.2\%$ when we consider the top $k=5$ tokens at each decoding step (Fig. 3). Further increasing $k$ to $10$ incurs a higher computational cost without raising the proportion of decoded golden tools.
A.2 Temperature Sampling
We experiment with temperature sampling and find that it is not an optimal strategy to gather candidate tools, as lower temperatures do not lead to enough diversity while higher ones generate irrelevant tools.
A.3 Tool Selection by Self-Consistency
We select tools by self-consistency on FuncQA’s validation set. We try choosing the tool that is most represented in the candidate set both before and after argument insertion. In both cases, We find that tool choice by self-consistency performs poorly (over 6% below ToolkenGPT in the best case).
Figure 3: Proportion of golden tools across the FuncQA validation set, for k=1, k=5 and k=10, where k is the number of top-probability tokens at each decoding step.
Appendix B Model Implementation and Training
@@
B.1 Tools
GSM8K-XL and FuncQA Hao et al. (2023) are annotated with four and 13 tools respectively, each representing a math operation. We illustrate these in Table 3. Tools are trained as additional tokens added to the standard language modelling head of Llama 3, which comprises 128,256 token representations. Therefore, we extend these representations to 128,269 in the case of FuncQA and 128,260 for GSM8K-XL.
B.2 Details of the Training Process
The LM head of Tecton consists of the standard head of Llama 3 8B Instruct Dubey et al. (2024) concatenated with an additional linear layer of size $embedding\_dimension\times number\_of\_tools$. The tool token embeddings are randomly initialized and trained on math reasoning QA pairs from GSM8K-XL and FuncQA, following a standard language modelling objective. Both datasets are annotated with the token positions at which each tool should be generated. Note that the original datasets made available by Hao et al. (2023) are annotated according to SentencePiece tokenization Kudo and Richardson (2018). We edit the annotations to be compatible with Llama 3’s Byte-Pair Encoding tokenizer Sennrich et al. (2016). We tune two distinct sets of special tokens, one on each dataset, as each requires a different set of tools as shown in Table 3. We train at half precision (FP16) for ten epochs, using learning rates {$1$e44-4- 4, $1$e33-3- 3}, batch size 1, and saving checkpoints at each epoch. We select the best checkpoint by measuring performance on a validation set.
@@ B.2 Details of the Training Process
Table 4: Sizes of our training and testing datasets. Note that we use ASDiv-XL, MAWPS-XL and SVAMP-XL only for testing. For FuncQA the reported test data size includes both splits—FuncQA-OH and FuncQA-MH.
@@
B.3 Inference-time Hyperparameters